# Graph Neural Preconditioners for Iterative Solutions of Sparse Linear Systems

### v1

@@ 2.1 Flexible GMRES
Algorithm 1 summarizes the solver. The algorithm assumes a restart length $m$. Starting from an initial guess $\mathbf{x}_{0}$ with residual vector $\mathbf{r}_{0}$ and 2-norm $\beta$, FGMRES runs $m$ Arnoldi steps, resulting in the relation
We provide a convergence analysis below, which is probably novel.
Theorem 1.
Let us consider the two terms on the right-hand side of (2). The first term $\kappa_{2}(\mathbf{X})\epsilon^{(m)}(\bm{\Lambda})\|\mathbf{r}_{0}\|_{2}$ is the standard convergence result of GMRES on the matrix $\mathbf{A}\widetilde{\mathbf{M}}$. The factor $\epsilon^{(m)}(\bm{\Lambda})$ is approximately exponential in $m$, when the eigenvalues $\lambda_{i}$ are located in an ellipse which excludes the origin (see [31, Corollary 6.33]). Furthermore, when the eigenvalues are closer to each other than to the origin, the base of the exponential is close to zero, resulting in rapid convergence. Meanwhile, when $\mathbf{A}\widetilde{\mathbf{M}}$ is close to normal, $\kappa_{2}(\mathbf{X})$ is close to 1.
What the nonlinear operator preconditioner $\mathbf{M}$ affects is the second term $\|\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}-\widetilde{\mathbf{Q}}_{m}\widetilde{\mathbf{Q}}_{m}^{\top}\|_{2}\|\mathbf{r}_{0}\|_{2}$. This term does not show convergence on the surface, but it offers some intuition. Assuming no breakdown, one can always find an $\widetilde{\mathbf{M}}$ such that this term vanishes. Indeed, $\widetilde{\mathbf{M}}=\mathbf{A}^{-1}\mathbf{V}_{n}\mathbf{H}_{n}\mathbf{V}_{n}^{\top}$ makes $\mathbf{A}\mathbf{Z}_{m}=\mathbf{A}\widetilde{\mathbf{Z}}_{m}$, and hence $\mathbf{Q}_{m}=\widetilde{\mathbf{Q}}_{m}$, for all $m$. When the preconditioner $\mathbf{M}$ is close to $\mathbf{A}^{-1}$, $\mathbf{A}\mathbf{Z}_{m}$ is a small perturbation of $\mathbf{V}_{m}$ by definition. Then, (1) suggests that $\overline{\mathbf{H}}_{m}$ is close to the identity matrix for all $m$ and so is $\mathbf{H}_{n}$. Thus, the following corollary characterizes the convergence of the residual.
Corollary 2.
2.2 Training Data Generation
We generate a set of $(\mathbf{b},\mathbf{x})$ pairs to train $\mathbf{M}:\mathbf{b}\to\mathbf{x}$. Unlike neural operators where the data generation is costly (e.g., requiring solving PDEs), which limits the training set size, for linear systems, we can sample $\mathbf{x}$ from a certain distribution and obtain $\mathbf{b}=\mathbf{A}\mathbf{x}$ with negligible cost, which creates a streaming training set of $(\mathbf{b},\mathbf{x})$ pairs without size limit.
There is no free lunch. One may sample $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$, which leads to $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{A}\mathbf{A}^{\top})$. This distribution is skewed toward the dominant eigen-subspace of $\mathbf{A}\mathbf{A}^{\top}$. Hence, using samples from it for training may result in a poor performance of the preconditioner, when applied to inputs lying close to the bottom eigen-subspace. One may alternatively want the training data $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$, which covers uniformly all possibilities of the preconditioner input on a sphere. However, there are two challenges. First, it requires sampling $\mathbf{x}$ from $\mathcal{N}(\mathbf{0},\mathbf{A}^{-1}\mathbf{A}^{-\top})$, which is a chicken-and-egg problem. Second, $\mathcal{N}(\mathbf{0},\mathbf{A}^{-1}\mathbf{A}^{-\top})$ is also skewed in the spectrum, causing difficulty in network training.
Then, with simple algebra we obtain that
The covariance of $\mathbf{b}$, $\bm{\Sigma}_{m}^{\mathbf{b}}$, is a projector, because both $\mathbf{V}_{m+1}$ and $\mathbf{W}_{m}$ are orthonormal. As $m$ grows, this covariance tends to the identity $\mathbf{I}_{n}$. By the theory of the Arnoldi process [24], $\mathbf{V}_{m+1}$ contains better and better approximations of the extreme eigenvectors as $m$ increases. Hence, $\mathcal{N}(\mathbf{0},\bm{\Sigma}_{m}^{\mathbf{b}})$ has a high density close to the bottom eigen-subspace of $\mathbf{A}$. Of course, because $\bm{\Sigma}_{m}^{\mathbf{b}}$ is low-rank, the distribution has zero density along many directions. Therefore, we sample $\mathbf{x}$ from both $\mathcal{N}(\mathbf{0},\bm{\Sigma}_{m}^{\mathbf{x}})$ and $\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ to form each training batch, to cover more important subspace directions.
2.3 Scale-Equivariant Graph Neural Preconditioner
We now define the neural network $\mathbf{M}$ that approximates $\mathbf{A}^{-1}$. Naturally, a (sparse) matrix $\mathbf{A}=[a_{ij}]$ admits a graph interpretation, similar to how AMG interpreting the coefficient matrix with a graph. Treating $\mathbf{A}$ as the graph adjacency matrix allows us to use GNNs [16, 10, 32, 34] to parameterize the preconditioner. These neural networks perform graph convolutions in a neighborhood of each node, reducing the quadratic pairwise cost by ignoring nodes faraway from the neighborhood.
Our neural network is based on the graph convolutional network (GCN) [16]. A GCN layer is defined as $\text{GCONV}(\mathbf{X})=\operatorname{ReLU}(\widehat{\mathbf{A}}\mathbf{X}\mathbf{W})$, where $\widehat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ is some normalization of $\mathbf{A}$, $\mathbf{X}\in\mathbb{R}^{n\times d_{\text{in}}}$ is input data, and $\mathbf{W}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is a learnable parameter.
  1. 1.
    We normalize $\mathbf{A}$ differently. Specifically, we define
    The factor $\gamma$ is an upper bound of the spectral radius of $\mathbf{A}$ based on the Gershgorin circle theorem [9]. This avoids division-by-zero that may occur in the standard GCN normalization, because the nonzeros of $\mathbf{A}$ can be both positive and negative.
  2. 2.
    We add a residual connection to GCONV, to allow stacking a deeper GNN. Specifically, we define a Res-GCONV layer to be
[…]
  3. 3.
Additionally, note that the operator to approximate, $\mathbf{A}^{-1}$, is scale-equivariant, but a neural network is not guaranteed so. More importantly, the neural network input $\mathbf{b}$ may have very different scales, due to the sampling approach considered in the preceding subsection. Hence, to facilitate training, we design a parameter-free scaling $s$ and back-scaling $s^{-1}$:
3 Evaluation Methodology and Experiment Setting
Problems. Our evaluation strives to cover a large number of problems, which come from as diverse application areas as possible. To this end, we turn to the SuiteSparse matrix collection https://sparse.tamu.edu, which is a widely used benchmark in numerical linear algebra. We select all square, real-valued, and non-SPD matrices whose number of rows falls between 1K and 100K and whose number of nonzeros is fewer than 2M. This selection results in 867 matrices with varying densities and applications. Some applications are involved with PDEs (such as computational fluid dynamics), while others come from graph, optimization, economics, and statistics problems.
To facilitate solving the linear systems, we prescale each $\mathbf{A}$ by $\gamma$ defined in (6). This scaling is a form of normalization, which avoids the overall entries of the matrix being exceedingly small or large. All experiments assume the ground truth solution $\mathbf{x}=\mathbf{1}$. FGMRES starts with $\mathbf{x}_{0}=\mathbf{0}$.
Compared methods. We compare GNP with three widely used general-purpose preconditioners: ILU, AMG, and GMRES (using GMRES to precondition GMRES is also called inner-outer GMRES). The endeavor of preconditioning over 800 matrices from different domains limits the choices, but these preconditioners are handy with Python software support. ILU comes from scipy.sparse.linalg.spilu; it in turn comes from SuperLU [18, 19], which implements the thresholded version of ILU [28]. We use the default drop tolerance and fill factor without tuning. AMG comes from PyAMG [1]. Specifically, we use pyamg.blackbox.solver().aspreconditioner as the preconditioner, with the configuration of the blackbox solver computed from pyamg.blackbox.solver_configuration. GMRES is self-implemented. For (the inner) GMRES, we use 10 iterations and stop it when it reaches a relative residual norm tolerance of 1e-6. For (the outer) FGMRES, the restart cycle is 10.
Stopping criteria. The stopping criteria of all linear system solutions are rtol = 1e-8 and maxiters = 100.
Evaluation metrics. Evaluating the solution qualities and comparing different methods for a large number of problems require programmable metrics, but defining these metrics is harder than it appears to be. There are several reasons: (1) not every method reaches rtol within maxiters; (2) the iteration cost for each method is different; and (3) the convergence curves of different methods may cross over. Hence, comparing the number of iterations, the final relative residual norm, or the time, alone, is not enough. We propose two novel metrics. The first one, area under the relative residual norm curve with respect to iterations, is defined as
[…]
where iters is the actual number of iterations when FGMRES stops and $r_{i}$ is the relative residual norm at iteration $i$. We take logarithm because residual plots are typically done in the log scale. This metric compares methods based on the iteration count, taking into account the history rather than the final error alone.
[…]
where $T$ is the elapsed time when FGMRES stops, $r(t)$ is the relative residual norm at time $t$, and $t_{i}$ is the elapsed time at iteration $i$. This metric compares methods based on the solution time, taking into account the history of the errors.
Hyperparameters. To feasibly train over 800 GNNs, we use the same set of hyperparameters for each of them without tuning. There is a strong potential that the GNN performance can be greatly improved with careful hyperparameter tuning, or even problem-specific tuning. Our purpose is to understand the general behavior of GNP, particularly its strengths and weaknesses. We use $L=8$ Res-GCONV layers, set the layer input/output dimension to 16, and use 2-layer MLPs with hidden dimension 32 for lifting/projection. We use Adam [15] as the optimizer, set the learning rate to 1e-3, and train for 2000 steps with a batch size of 16. We apply neither dropouts nor weight decays. Because data are sampled from the same distribution, we use the model at the best training epoch as the preconditioner, without resorting to a validation set.
We use the $\ell_{1}$ residual norm $\|\mathbf{A}\mathbf{M}(\mathbf{b})-\mathbf{A}\mathbf{x}\|_{1}$ as the training loss. We use $m=40$ Arnoldi steps when sampling the $(\mathbf{x},\mathbf{b})$ pairs according to (5). Among the 16 pairs in a batch, 8 pairs follow (5) and 8 pairs follow $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$.
Compute environment. Our experiments are conducted on a machine with one Tesla V100(16GB) GPU, 96 Intel Xeon 2.40GHz cores, and 386GB main memory. All code is implemented in Python with Pytorch.
4 Results
We organize the experiment results by key questions of interest.
How does GNP perform compared with traditional preconditioners? We show in Figure 2 the percentage of problems for which each preconditioner performs the best, with respect to iteration counts and solution time (see the Iter-AUC and Time-AUC metrics defined in the preceding section). GNP performs the best for a substantial portion of the problems. In comparison, ILU and AMG are more competitive, but they are inferior in some aspects, which will be revealed bit by bit later. GMRES as a preconditioner is also competitive in iteration counts, but less so in solution time compared with GNP. Interestingly, for nearly half of the problems, not using a preconditioner works the best in the time metric. However, in most of these problems the attainable residual norm is the worst if no preconditioner is used.
Figure 2: Percentage of problems for which each preconditioner performs the best.
For what problems does GNP perform the best? We expand the granularity of Figure 2 into Figure 3, overlaid with the size distribution and the condition-number distribution of the problems. A key finding to note is that GNP is particularly useful for ill-conditioned matrices (i.e., those with a condition number $\geq 10^{16}$). ILU is less competitive in these problems, compared with other preconditioners, possibly because it focuses on the sparsity structure, whose connection with the spectrum is less clear.
On the other hand, the matrix size does not affect much the GNP performance. One sees this more clearly when consulting Figure 7 in the appendix, which is a “proportion” version of Figure 3: the proportion of problems for which GNP performs the best is relatively flat across matrix sizes. Similarly, symmetry also does not play a distinguishing role: GNP performs the best for 8.8% of the symmetric matrices and 17.7% of the nonsymmetric matrices.
Figure 3: Breakdown of best preconditioners with respect to matrix sizes and condition numbers. The last bar on the right plot is for condition number $\geq 10^{16}$.
What does the convergence history look like with GNP? In Figure 4, we show a few examples when GNP performs the best. It is worthwhile to note that in this case, ILU fails for most of the problems, either in the construction phase or in the solution phase (more on this later). These examples indicate that the convergence of GNP either tracks that of other competitive preconditioners with a similar rate (see problems Simon/venkat25 and VDOL/kineticBatchReactor_5), or becomes significantly faster (see problems GHS_indef/bratu3d and VanVelzen/std1_Jac3). Notably, on VanVelzen/std1_Jac3, GNP uses only four iterations to reach 1e-8, while other preconditioners take 100 iterations but still cannot decrease the relative residual norm to below 1e-2.
Figure 4: Example convergence histories.
We additionally plot the convergence curves with respect to time in Figure 8 in the appendix. An important finding to note is that GMRES as a preconditioner generally takes much longer time to run than GNP. This is because GMRES is bottlenecked by the orthogonalization process, shadowing the trickier cost comparison between more matrix-vector multiplications in GMRES and fewer matrix-matrix multiplications in GNP, with respect to the matrix $\mathbf{A}$.
How does the construction cost of GNP fare, compared with that of other preconditioners? GMRES does not require construction; hence, we mainly compare GNP with ILU and AMG. In Figure 5, we plot their construction times. The training time of GNP is nearly proportional to the matrix size, as expected. In contrast, the construction time of ILU and AMG is much less predictable. Even though for many of the problems, the time is only a fraction of that of GNP, there exist quite a few cases where the time is significantly longer. In the worst case, the construction of the AMG preconditioner is more than an order of magnitude more costly than that of GNP. Moreover, it is unclear when such “outliers” happen, as they do not seem to be correlated with the matrix size.
Figure 5: Preconditioner construction time.
How robust is GNP? We summarize in Table 1 the number of failures for each preconditioner. GNP and GMRES are highly robust. We see that neural network training does not cause any troubles, which is a practical advantage. In contrast, ILU fails for nearly half of the problems and AMG fails for nearly 8%. According to the error log, the common failures of ILU construction are that “(f)actor is exactly singular” and that “matrix is singular … in file ilu_dpivotL.c”, while the common failures of AMG are “array … contain(s) infs or NaNs”. Meanwhile, solution failures occur when the residual norm tracked by the QR factorization of the upper Hessenberg $\overline{\mathbf{H}}_{m}$ fails to match the actual residual norm. While these two preconditioners perform the best for the most number of problems, safeguarding their robustness on other problems remains challenging.
Table 1: Failures of preconditioners.
GNP ILU AMG GMRES
Construction failure 0 (0.00%) 348 (40.14%) 62 (7.15%) N/A
Solution failure 1 (0.12%) 61 (07.04%) 5 (0.58%) 1 (0.12%)
Are the proposed training data generation and the scale-equivariance design necessary? In Figure 6, we compare the proposed designs versus alternatives, for both the training loss and the relative residual norm. In the comparison of training data generation, using $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ leads to lower losses, while using $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ leads to higher losses. This is expected, because the former expects similar outputs from the neural network, which is easier to train, while the latter expects drastically different outputs and makes the network difficult to train. Using $\mathbf{x}$ from the proposed mixture leads to losses lying in between. This case leads to the best preconditioning performance—more problems having lower residual norms (especially those near rtol). In other words, the proposed training data generation strikes the best balance between easiness of neural network training and goodness of the resulting preconditioner.
5 Discussions and Conclusions
We have presented a GNN approach to preconditioning Krylov solvers for the solution of large, sparse linear systems. This is the first work in our knowledge to use a neural network as an approximation to the matrix inverse. Compared with traditional preconditioners, this approach is robust with predictable construction costs and is effective for highly ill-conditioned problems. We evaluate the approach on more than 800 matrices from over 20 application areas, the widest coverage in the literature.
The work can be extended in many avenues. First, an immediate follow up is the preconditioning of SPD matrices. While flexible CG was analyzed in the past [25], it is relatively fragile with respect to a large variation of the preconditioner. We speculate that a split preconditioner can work more robustly and some form of autoencoder networks better serves this purpose.
Second, while GNP is trained for an individual matrix in this work for simplicity, it is possible to extend it to a sequence of evolving matrices, such as in sequential problems or time stepping. Continual learning, which is concerned with continuously adapting a trained neural network for slowly evolving data distributions, can amortize the initial preconditioner construction cost with more and more linear systems in sequel.
Third, the GPU memory limits the size of the matrices and the complexity of the neural networks that we can experiment with. Future work can explore multi-GPU and/or distributed training of GNNs [14, 13] for scaling GNP to even larger matrices. The training for large graphs typically uses neighborhood sampling to mitigate the unique “neighborhood explosion” problem in GNNs. Some sampling approaches, such as layer-wise sampling [5], are tied to sketching the matrix product $\widehat{\mathbf{A}}\mathbf{X}$ inside the graph convolution layers (7) with certain theoretical guarantees [4].
Appendix A Proofs
@@
see, e.g., [31, Proposition 6.32]. On the other hand, because $\mathbf{r}_{m}=\mathbf{r}_{0}-\mathbf{A}\mathbf{Z}_{m}\mathbf{y}_{m}$ where $\mathbf{y}_{m}$ minimizes $\|\mathbf{r}_{0}-\mathbf{A}\mathbf{Z}_{m}\mathbf{y}\|_{2}$, we have $\mathbf{r}_{m}=\mathbf{r}_{0}-(\mathbf{A}\mathbf{Z}_{m})(\mathbf{A}\mathbf{Z}_{m})^{+}\mathbf{r}_{0}=\mathbf{r}_{0}-\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}\mathbf{r}_{0}$. Therefore, $\|\mathbf{r}_{m}-\widetilde{\mathbf{r}}_{m}\|_{2}=\|\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}\mathbf{r}_{0}-\widetilde{\mathbf{Q}}_{m}\widetilde{\mathbf{Q}}_{m}^{\top}\mathbf{r}_{0}\|_{2}\leq\|\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}-\widetilde{\mathbf{Q}}_{m}\widetilde{\mathbf{Q}}_{m}^{\top}\|_{2}\|\mathbf{r}_{0}\|_{2}$. ∎
@@
Appendix B Additional Experiment Results

### v3

@@ Appendix B FGMRES
Appendix C Proofs
@@
see, e.g., Saad (2003, Proposition 6.32). On the other hand, because $\mathbf{r}_{m}=\mathbf{r}_{0}-\mathbf{A}\mathbf{Z}_{m}\mathbf{y}_{m}$ where $\mathbf{y}_{m}$ minimizes $\|\mathbf{r}_{0}-\mathbf{A}\mathbf{Z}_{m}\mathbf{y}\|_{2}$, we have $\mathbf{r}_{m}=\mathbf{r}_{0}-(\mathbf{A}\mathbf{Z}_{m})(\mathbf{A}\mathbf{Z}_{m})^{+}\mathbf{r}_{0}=\mathbf{r}_{0}-\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}\mathbf{r}_{0}$. Therefore, $\|\mathbf{r}_{m}-\widetilde{\mathbf{r}}_{m}\|_{2}=\|\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}\mathbf{r}_{0}-\widetilde{\mathbf{Q}}_{m}\widetilde{\mathbf{Q}}_{m}^{\top}\mathbf{r}_{0}\|_{2}\leq\|\mathbf{Q}_{m}\mathbf{Q}_{m}^{\top}-\widetilde{\mathbf{Q}}_{m}\widetilde{\mathbf{Q}}_{m}^{\top}\|_{2}\|\mathbf{r}_{0}\|_{2}$. ∎
@@
Appendix D Approximately Exponential Convergence of $\|\mathbf{r}_{m}\|_{2}$
We could bound the minimax polynomial $\epsilon^{(m)}$ by using Chebyshev polynomials to obtain an (approximately) exponential convergence of the residual norm $\|\mathbf{r}_{m}\|_{2}$ for large $m$.
Assume that all the eigenvalues $\sigma_{i}$ of $\bm{\Sigma}$ are enclosed by an ellipse $E(c,d,a)$ which excludes the origin, where $c$ is the center, $d$ is the focal distance, and $a$ is major semi-axis. We have
$\epsilon^{(m)}(\bm{\Sigma})\leq\min_{p\in\mathbb{P}_{m},\,p(0)=1}\max_{\sigma\in E(c,d,a)}|p(\sigma)|,$
by the fact that the maximum modulus of a complex analytical function is reached on the boundary of the domain. Then, we invoke Eqn (6.119) of Saad (2003) and obtain
$\epsilon^{(m)}(\bm{\Sigma})\leq\frac{C_{m}(a/d)}{C_{m}(c/d)},$
where $C_{m}$ denotes the (complex) Chebyshev polynomial of degree $m$. When $m$ is large,
$\frac{C_{m}(a/d)}{C_{m}(c/d)}\approx\left(\frac{a+\sqrt{a^{2}-d^{2}}}{c+\sqrt{c^{2}-d^{2}}}\right)^{m},$
which gives an (approximately) exponential function in $m$. Applying Corollary 2, we conclude that when $m$ is large,
$\|\mathbf{r}_{m}\|_{2}\lessapprox\kappa_{2}(\mathbf{Y})\|\mathbf{r}_{0}\|_{2}\left(\frac{a+\sqrt{a^{2}-d^{2}}}{c+\sqrt{c^{2}-d^{2}}}\right)^{m}.$
Appendix E Scalability
Based on the architecture outlined in Figure 1, the computational cost of GNP is dominated by matrix-matrix multiplications, where the left matrix is either $\mathbf{A}$ (e.g., in GCONV) or a tall one that has $n$ rows (e.g., in the MLP and Lin layers and in GCONV). This forward analysis holds true for training and inference, because the back propagation also uses these matrix-matrix multiplications. Hence, the computational cost of GNP scales as $O(\operatorname{nz}(\mathbf{A}))$, the same as that of the Krylov solver.
Practical considerations, on the other hand, are more sensitive to the constants hidden inside the big-O notation. For example, one needs to store many (but still a constant number of) vectors of length $n$, due to the batch size, the hidden dimensions of the neural network, and the automatic differentiation. Denote by $c$ this constant number; then, the GPU faces a storage pressure of $\operatorname{nz}(\mathbf{A})+cn$. Once this amount is beyond the memory capacity of the GPU, one either performs training (and even inference) by using CPU only, offloads some data to the CPU and transfers them back to the GPU on demand, or undertakes more laborious engineering by distributing $\mathbf{A}$ and other data across multiple GPUs. For the last option (multi-GPU and/or distributed training), the literature of GNN training brings in additional techniques (such as sampling and mini-batching the graph) to accelerate the computation (Kaler et al., 2022; 2023).
Appendix F Additional Experiment Results
F.1 Additional Metrics for Comparing Preconditioners
In addition to the Iter-AUC and Time-AUC metrics, one may be interested in using the final relative residual norm to compare the preconditioners. To this end, we summarize the results in Figure 7 by using this residual metric, as an extension of Figure 2. Note that for each problem, we solve the linear system twice, once using rtol and maxiters as the stopping criteria; and the other time using rtol and timeout. Hence, there are two charts in Figure 7, depending on how the iterations are terminated.
Figure 7: Percentage of problems on which each preconditioner performs the best.
We see that across these two termination scenarios, observations are similar. Compared to the use of the Iter-AUC and Time-AUC metrics, GNP is the best for a similar portion of problems under the residual metric. The portion of problems that ILU and GMRES (as a preconditioner) perform the best increases under this metric, while the portion that AMG performs the best decreases. Jacobi and no preconditioner remain less competitive.
F.2 Preconditioner Construction Time and Solution Time
In Figure 8, we plot the preconditioner construction time and solution time for each matrix and each preconditioner. This figure is a counterpart of Figure 3 by removing the log scale in the y axis. One sees that the GNP construction time is nearly proportional to the matrix size.
Figure 8: Preconditioner construction time and solution time (using maxiters to stop). The y axis is in the linear scale.
F.3 Breakdown of Best Preconditioners With Respect to Matrix Distributions